Fill in the blanks:
A) 1 lakh =________ ten thousand.
B) 1 million =________ hundred thousand.
C) 1 crore =________ ten lakh.
D) 1 crore =________ million.
E) 1 million =________ lakh.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
In this question it is given that we have to fill the blanks. To find the solution, we have to know how to represent the given quantity numerically and after that according to the question we have to transform them to the new quantity. Complete step by step answer: As we know that 1 lakh = 100000 (5 zeroes), so this can be written as, A) 1 lakh = 100000 = 10×10000 = 10 × ten thousand. Therefore, 1 lakh = 10 ten thousand. B) 1 million = 10,00000 (6 zeroes) = 10 Lakh and by previous we can write that, 10 lakh = 10 × ten thousand = 10 × 10 × 1000 = 10 × 10 × 1 thousand = Hundred thousand. So, 1 million can be written as 10 × hundred thousand. C) Now 1 crore =100 Lakh, which can be written as, 1 crore = 10 ×10 Lakh=10 ten Lakh. Thus 1 crore = 10 ten lakh. D) Again, 1 crore = 100 lakh = 10 × 10 lakh = 10 × 1 million = 10 million. So, one crore is equal to ten million. E) And lastly, 1 million which can be written as 10 Lakh as seen above in (B). Therefore, required solution is A) 1 lakh = 10 ten thousand. B) 1 million = 10 hundred thousand. C) 1 crore = 10 ten lakh. D) 1 crore = 10 million. E) 1 million = 10 lakh. Note: While solving any number related problems you need to know that whenever we write zero alongside 1 then this will give us a new quantity. Like if we write 10 then this is called as ten. 100=1 hundred , 1000=1 thousand, 10000= ten thousand, 100000= 1lakh, 1000000= 10 lakh, 10000000=1 crore.